OpenStudy (anonymous):

help , I need you to show me the work for -3x+4<=x+8

OpenStudy (anonymous):

if the inequality sign was not there, and there was just an equal sign, could you solve this problem ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-3x + 4 <= x + 8 what you want to do is get all the x terms on one side of the inequality sign and everything else on the other side. But remember, what you do to one side, you have to do to the other side. -3x + 4 <= x + 8 -- first we will subtract 4 from both sides -3x + 4 - 4 <= x + 8 - 4 -- by doing this, we cancelled out the 4's on the left side. now we will combine like terms.. -3x <= x + 4 -- now we want to subtract x from both sides -3x - x <= x - x + 4 -- we have now cancelled the x's on the right side combine like terms -4x <= 4 -- now we have to divide by -4, one rule you have to remember with inequalities is that when dividing/multiplying by a negative number, it changes the inequality sign. -4x/-4 >= 4/-4 x > = -1 any questions ?

OpenStudy (anonymous):

thank you so much ! this helped a lot

OpenStudy (anonymous):

glad to help :) here is a little tip...when you want to move something across the inequality sign, you can do so by just changing the sign. 2x + 4 < 6.....you want to move the 4 to the other side, just move it and change the sign. 2x < 6 - 4....just a little tip you might find useful